OpenStudy (anonymous):

What is 21 divided by 3/7?

OpenStudy (anonymous):

21 divided by 3/7 is the same as 21 times 7/3. \[\frac{ 21 }{ 1 } \times \frac{ 7 }{ 3 }\]

OpenStudy (anonymous):

When multiplying fractions, multiply across.

OpenStudy (anonymous):

okay so i got 147/3?

OpenStudy (anonymous):

yes. Your teacher may want you to simplify it though. Here's a trick: